AN EFFORT FORCE OF 200.0 N IS APPLIED TO AN IDEAL MACHINE TO MOVE A 750.0-N RESISTANCE A DISTANCE OF 300.0 CM
A. THROUGH WHAT DISTANCE MUST THE EFFORT FORCE ACT
B.WHAT IS THE IMA OF THE MACHINE)
1
Expert's answer
2020-04-20T19:02:36-0400

As per the given question,

Force =200N

Resistance force =750N

Distance =300cm

a) Let the distance of the effort force act =d

Now, applying the conservation of energy,

"200\times d=750\times 300"

"\Rightarrow d=\dfrac{750\times 300}{200}=1125cn"

b) IMA="\dfrac{friction }{Force}=\dfrac{750}{200}=3.75"